### v2

# REFRAG: Rethinking RAG based Decoding

Abstract
Large Language Models (LLMs) have demonstrated remarkable capabilities in leveraging extensive external knowledge to enhance responses in multi-turn and agentic applications, such as retrieval-augmented generation (RAG). However, processing long-context inputs introduces significant system latency and demands substantial memory for the key-value cache, resulting in reduced throughput and a fundamental trade-off between knowledge enrichment and system efficiency. While minimizing latency for long-context inputs is a primary objective for LLMs, we contend that RAG systems require specialized consideration. In RAG, much of the LLM context consists of concatenated passages from retrieval, with only a small subset directly relevant to the query. These passages often exhibit low semantic similarity due to diversity or deduplication during re-ranking, leading to block-diagonal attention patterns that differ from those in standard LLM generation tasks. Based on this observation, we argue that most computations over the RAG context during decoding are unnecessary and can be eliminated with minimal impact on performance. To this end, we propose REFRAG, an efficient decoding framework that compresses, senses, and expands to improve latency in RAG applications. By exploiting this attention sparsity structure, we demonstrate a $30.85\times$ the time-to-first-token acceleration ($3.75\times$ improvement to previous work) without loss in perplexity. In addition, our optimization framework for large context enables REFRAG to extend the context size of LLMs by $16\times$. We provide rigorous validation of REFRAG across diverse long-context tasks, including RAG, multi-turn conversations, and long document summarization, spanning a wide range of datasets. Experimental results confirm that REFRAG delivers substantial speedup with no loss in accuracy compared to LLaMA models and other state-of-the-art baselines across various context sizes. Additionally, our experiments establish that the expanded context window of REFRAG further enhances accuracy for popular applications.
[…]
1 Introduction
Large Language Models (LLMs) have demonstrated impressive capabilities in contextual learning, leveraging information from their input to achieve superior performance across a range of downstream applications. For instance, in multi-turn conversations (Roller et al., 2021; Zhang et al., 2020), incorporating historical dialogue into the context enables LLMs to respond more effectively to user queries. In retrieval-augmented generation (RAG) (Guu et al., 2020; Izacard et al., 2022), LLMs generate more accurate answers by utilizing relevant search results retrieved from external sources. These examples highlight the power of LLMs to learn from context. However, it is well established that increasing prompt length for contextual learning leads to higher latency and greater memory consumption during inference (Yen et al., 2024). Specifically, longer prompts require additional memory for the key-value (KV) cache, which scales linearly with prompt length. Moreover, the time-to-first-token (TTFT) latency increases quadratically, while the time-to-iterative-token (TTIT) latency grows linearly with prompt length (Liu et al., 2025). As a result, LLM inference throughput degrades with larger contexts, limiting their applicability in scenarios demanding high throughput and low latency, such as web-scale discovery. Therefore, developing novel model architectures that optimize memory usage and inference latency is crucial for enhancing the practicality of contextual learning in these applications.
Optimizing inference latency for LLMs with extensive context is an active area of research, with approaches ranging from modifying the attention mechanism’s complexity (Beltagy et al., 2020) to sparsifying attention and context (Child et al., 2019; Xiao et al., 2024; Jiang et al., 2024), and altering context feeding strategies (Yen et al., 2024). However, most existing methods target generic LLM tasks with long context and are largely orthogonal to our work. This paper focuses on RAG-based applications, such as web-scale search, with the goal of improving inference latency, specifically, the TTFT. We argue that specialized techniques exploiting the unique structure and sparsity inherent in RAG contexts can substantially reduce memory and computational overhead. Treating RAG TTFT as a generic LLM inference problem overlooks several key aspects: 1) Inefficient Token Allocation. RAG contexts often contain sparse information, with many retrieved passages being uninformative and reused across multiple inferences. Allocating memory/computation for all the tokens, as we show in this paper, is unnecessarily wasteful. 2) Wasteful Encoding and Other Information. The retrieval process in RAG has already pre-processed the chunks of the contexts, and their encodings and other correlations with the query are already available due to the use of vectorizations and re-rankings. This information is discarded during decoding. 3) Unusually Structured and Sparse Attention. Due to diversity and other operations such as deduplication, most context chunks during decoding are unrelated, resulting in predominantly zero cross-attention between chunks (see figure˜7).
1.1 Our Contributions
We propose REFRAG (REpresentation For RAG), a novel mechanism for efficient decoding of contexts in RAG. REFRAG significantly reduces latency, TTFT, and memory usage during decoding, all without requiring modifications to the LLM architecture or introducing new decoder parameters.
REFRAG makes several novel modifications to the decoding process: Instead of using tokens from retrieved passages as input, REFRAG leverages pre-computed, compressed chunk embeddings as approximate representations, feeding these embeddings directly into the decoder. This approach offers three main advantages: 1) It shortens the decoder’s input length, improving token allocation efficiency; 2) It enables reuse of pre-computed chunk embeddings from retrieval, eliminating redundant computation; and 3) It reduces attention computation complexity, which now scales quadratically with the number of chunks rather than the number of tokens in the context. Unlike prior methods (Yen et al., 2024), REFRAG supports compression of token chunks at arbitrary positions (see figure˜1) while preserving the autoregressive nature of the decoder, thereby supporting multi-turn and agentic applications. This “compress anywhere” capability is further enhanced by a lightweight reinforcement learning (RL) policy that selectively determines when full chunk token input is necessary and when low-cost, approximate chunk embeddings suffice . As a result, REFRAG minimizes reliance on computationally intensive token embeddings, condensing most chunks for the query in RAG settings.
We provide rigorous experimental validations of the effectiveness of REFRAG in continual pre-training and many real word long-context applications including RAG, multi-turn conversation with RAG and long document summarization. Results show that we achieve $30.75\times$ TTFT acceleration without loss in perplexity which is $3.75\times$ than previous method. Moreover, with extended context due to our compression, REFRAG achieves better performance than LLaMA without incurring higher latency in the downstream applications.
2 Model Architecture
Figure 1: The main design of REFRAG. The input context is chunked and processed by the light-weight encoder to produce chunk embeddings, which are precomputable for efficient reuse. A light-weight RL policy decide few chunks to expand. These chunk embeddings along with the token embeddings of the question input are fed to the decoder.
We denote the decoder model as $\mathcal{M}_{\text{dec}}$ and the encoder model as $\mathcal{M}_{\text{enc}}$. Given an input with $T$ tokens $x_{1},x_{2},\dots,x_{T}$, we assume that the first $q$ tokens are main input tokens (e.g., questions) and the last $s$ tokens are context tokens (e.g., retrieved passages in RAG). We have $q+s=T$. For clarity, we focus on a single turn of question and retrieval in this section.
Model overview. Figure˜1 shows the main architecture of REFRAG. This model consists of a decoder-only foundation model (e.g., LLaMA (Touvron et al., 2023)) and a lightweight encoder model (e.g., Roberta (Liu et al., 2019)). When given a question $x_{1},\dots,x_{q}$ and context $x_{q+1},\dots,x_{T}$ and , the context is chunked into $L\coloneq\frac{s}{k}$ number of $k$-sized chunks $\{C_{1},\dots,C_{L}\}$ where $C_{i}=\{x_{q+k*i},\dots,x_{q+k*i+k-1}\}$. The encoder model then processes all the chunks to obtain a chunk embedding for each chunk ${\mathbf{c}}_{i}=\mathcal{M}_{\text{enc}}(C_{i})$. This chunk embedding is then projected with a projection layer $\phi$ to match the size of the token embedding of the decoder model, ${\mathbf{e}}^{\text{cnk}}_{i}=\phi({\mathbf{c}}_{i})$. These projected chunk embeddings are then fed to the decoder model along with the token embeddings for the question to generate the answer $y\sim\mathcal{M}_{\text{dec}}(\{{\mathbf{e}}_{1},\dots,{\mathbf{e}}_{q},{\mathbf{e}}^{\text{cnk}}_{1},\dots,{\mathbf{e}}^{\text{cnk}}_{L}\})$ where ${\mathbf{e}}_{i}$ is the token embedding for token $x_{i}$. In real applications (e.g., RAG), the context is the dominating part of the input (i.e., $s\gg q$) and hence the overall input to the decoder will be reduced by a factor of $\simeq k$. This architectural design leads to significant reductions in both latency and memory usage, primarily due to the shortened input sequence. Additionally, an RL policy is trained to do selective compression to further improve the performance which we will defer the discussion to section˜2. Next, we analyze the system performance gains achieved with a compression rate of $k$.
Figure 2: Empirical verification of inference acceleration of REFRAG with $k=16$.
Latency and throughput improvement. We evaluate three metrics: TTFT, the latency to generate the first token; TTIT, the time to generate each subsequent token; and Throughput, the number of tokens generated per unit time. Theoretical analysis (section˜9) shows that for short context lengths, our method achieves up to $k\times$ acceleration in TTFT and throughput. For longer context length, acceleration reaches up to $k^{2}\times$ for both metrics. Empirically, as shown in figure˜2, with a context length of $16384$ (mid-to-long context), REFRAG with $k=16$ achieves $16.53\times$ TTFT acceleration with cache and $8.59\times$ without cache [^1] REFRAG without cache means that we recompute the chunk embedding for the context and take this latency into account., both surpassing CEPE ($2.01\times$ and $1.04\times$, respectively), while achieving 9.3% performance (measured by log-perplexity) compared to CEPE (table˜1). We achieve up to $6.78\times$ throughput acceleration compared to LLaMA, significantly outperforming CEPE. With $k=32$, TTFT acceleration reaches $32.99\times$ compared to LLaMA ($3.75\times$ compared to CEPE) while maintaining similar performance to CEPE (see figure˜8 and table˜2). More detailed discussion on empirical evaluation is in section˜9.
3 Methodology
To align the encoder and decoder, we follow the work of Yen et al. (2024) to use the next paragraph prediction tasks for continual pre-training (CPT). Specifically, for each data data point, it contains $s+o=T$ number of tokens, which we use for CPT to prepare the model for downstream tasks utilizing chunk embeddings. To further enhance performance, we introduce selective compression via RL. After aligning the encoder and decoder through CPT, we apply supervised fine-tuning (SFT) to adapt the model to specific downstream tasks, such as RAG and multi-turn conversation. Additional details are provided in section˜5.
During CPT, we input the first $s$ tokens $x_{1:s}$ into the encoder and use its output to assist the decoder in predicting the next $o$ tokens $x_{s+1:s+o}$. This task encourages the model to leverage contextual information for next-paragraph prediction, thereby equipping it for downstream applications. The objective is to align any encoder–decoder combination so that the generations produced with compressed context closely resemble those generated by the original decoder with access to the full context.
3.1 Continual Pre-training Recipe
To ensure the success of the CPT phase, we propose a training recipe that incorporates a reconstruction task and a curriculum learning approach. Ablation studies in section˜4 demonstrate that this recipe is crucial for achieving strong CPT performance.
Reconstruction task. We input the first $s$ tokens $x_{1:s}$ to the encoder and learn to reconstruct tokens $x_{1:s}$ in the decoder. In this task, we freeze the decoder model and only train the encoder and projection layer. The main objectives are to align the encoder and projection layer so that: 1) encoder can compress $k$ tokens with minimal information loss, and 2) projection layer can effectively map the encoder’s chunk embeddings into the decoder’s token space, allowing the decoder to interpret and accurately reconstruct the original information. The reconstruction task was specifically chosen to encourage the model to rely on context memory rather than its parametric memory during training. Once the encoder is aligned with the decoder through this reconstruction task, we initiate CPT by unfreezing the decoder.
Curriculum learning. The training tasks described in the previous section may seem straightforward, but they are inherently complex. As the chunk length $k$ increases, the number of possible token combinations expands exponentially, specifically at a rate of $V^{k}$, where $V$ is the vocabulary size. Effectively capturing this diversity within a fixed-length embedding presents a significant challenge. Additionally, reconstructing $s=k\times L$ tokens from $L$ chunk embeddings further compounds the difficulty of the task.
Counterintuitively, directly continuing pre-training of the decoder to utilize encoder outputs did not reduce perplexity, even for the reconstruction task. To address the optimization challenge, we propose employing curriculum learning for both tasks. Curriculum learning incrementally increases task difficulty, enabling the model to gradually and effectively acquire complex skills. For the reconstruction task, training begins with reconstructing a single chunk: the encoder receives one chunk embedding ${\mathbf{c}}_{1}$ for $x_{1:k}$ and and the decoder reconstructs the $k$ tokens using the projected chunk embedding ${\mathbf{e}}^{\text{cnk}}_{1}$. Subsequently, the model reconstructs $x_{1:2k}$ from ${\mathbf{e}}^{\text{cnk}}_{1},{\mathbf{e}}^{\text{cnk}}_{2}$, and so forth. To continuously adjust task difficulty, we vary the data mixture over time, starting with examples dominated by easier tasks (e.g., single chunk embedding) and gradually shifting towards those dominated by more difficult tasks (i.e., $L$ chunk embeddings). A visualization of the data mixture during curriculum learning is provided in figure˜6, and the detailed scheduling is presented in table˜8.
Selective compression REFRAG introduces selective token compression, expanding important context chunks uncompressed to improve answer prediction. A RL policy, guided by next-paragraph prediction perplexity as a negative reward, determines which chunks to retain in their original form. The encoder and decoder are fine-tuned to handle mixed inputs of compressed and uncompressed chunks. The policy network leverages chunk embeddings and masking to optimize sequential chunk expansion, thereby preserving the decoder’s autoregressive property and enabling flexible placement of compression. Further discussion on sequential selection is provided in section˜9.1.
4 Experimental Results
Training datasets. We use the Slimpajama dataset (Soboleva et al., 2023), an open source dataset for LLM pre-training. This dataset contains data from Wikipedia, Arxiv, Books, StackExchange, GitHub, Commoncrawl, C4. We only use the Book and ArXiv domains from the dataset since these two domains contain long texts (Yen et al., 2024). We sampled from this dataset to construct a $20\text{B}$ token training dataset which contains $50\%$ data from Arxiv and $50\%$ data from Book.
Evaluation datasets. We report the performance on the Book and ArXiv domain from Slimpajama which are hold out for evaluation only. To inspect the generalization of the model, we also report results on the PG19 (Rae et al., 2019) and Proof-pile datasets (Azerbayev et al., 2023).
[…]
Table˜1 reports performance for $s=2048$ and $o\in\{512,1024,2048\}$, where, e.g., P512 denotes $o=512$. Bolded results compare baselines, excluding LLaMA-Full Context and LLaMA-32K, which use full context without compression and are expected to perform best. Notably, $\textsc{REFRAG}_{8}$ and $\textsc{REFRAG}_{16}$ consistently outperform other baselines across nearly all settings, while also achieving lower latency than CEPE (figure˜2). For reference, ${\textsc{LLaMA}}_{256}$ uses only the last 256 tokens, matching the number of chunk embeddings in $\textsc{REFRAG}_{8}$ ($s/k=256$), yet $\textsc{REFRAG}_{8}$ consistently surpasses $\textsc{LLaMA}_{256}$, demonstrating the effectiveness of compressed chunk embeddings.
Table˜2 evaluates $o=2048$ with extended context lengths $s\in\{4096,8192,16384\}$. Although our model is trained on $s+o=6144$, both $\textsc{REFRAG}_{8}$ and $\textsc{REFRAG}_{16}$ maintain superior performance at longer contexts. The original Llama-2-7B supports only a 44k context window, whereas our approach enables extrapolation via chunk embeddings, extending context and supporting broader applications.
With a compression rate of $16$, we achieve a $9.3\%$ average log-perplexity improvement over CEPE across four datasets [^2] Percentage calculated as $\frac{\textsc{LLaMA-No Context}-\text{Log-perplexity to inspect}}{\textsc{LLaMA-No Context}-\min(\textsc{LLaMA-Full Context},\textsc{LLaMA-32K})}$. Meanwhile, our method is $16.53\times$ faster than LLaMA in TTFT and $2.01\times$ faster than CEPE (section˜10.4). At a compression rate of $32$, our log-perplexity matches CEPE, while TTFT acceleration increases to $30.85\times$ over LLaMA and $3.75\times$ over CEPE.
Figure˜3 presents the performance of various methods for selective compression. We expand $p$ fraction of the chunks in the original token space using the RL policy. The effective compression rate $\frac{k}{1-p+kp}$ decreases when fewer chunks are compressed (i.e., $p$ increases). We compare the perplexity of $x_{s+1:s+o}$ using different selection policy under different $p$. The perplexity-based selection is an heuristic based selection which compresses chunks with low perplexity (Perplexity-desc) or high perplexity (Perplexity-asc). The perplexity is measured by the LLaMA-2-7B model. Intuitively, a chunk with lower perplexity contains less information and can therefore be compressed with minimal information loss. Ideally, this approach should outperform random selection, which is indeed observed in figure˜3. The RL-based selective compression policy consistently achieves superior performance across varying compression rates $p$.
Table 1: Log-Perplexity on output tokens $x_{s+1:s+o}$ given context tokens $x_{1:s}$ for different models. We use $s=2048$ and $o\in\{512,1024,2048\}$ here. Bolding are based on comparing baselines excluding LLaMA-Full Context and LLaMA-32K since they are expected to be the best (ideally). The lower the better ($\mathbf{\downarrow}$).
Arxiv Book PG19 ProofPile
P512 P1024 P2048 P512 P1024 P2048 P512 P1024 P2048 P512 P1024 P2048 $\mathbf{\downarrow}$
LLaMA-Full Context 1.075 1.074 1.069 1.830 1.827 1.826 1.947 1.941 1.935 0.952 0.940 0.931
LLaMA-32K 1.086 1.084 1.076 1.887 1.883 1.880 1.988 1.982 1.975 0.961 0.948 0.937
LLaMA-No Context 1.526 1.371 1.254 2.101 1.995 1.927 2.211 2.102 2.030 1.437 1.256 1.127
$\textsc{LLaMA}_{256}$ 1.267 1.221 1.171 1.924 1.897 1.874 2.031 2.003 1.978 1.156 1.094 1.038
REPLUG 1.526 1.371 1.254 2.101 1.995 1.927 2.211 2.102 2.030 1.437 1.256 1.127
CEPE 1.196 1.148 1.107 1.946 1.896 1.864 2.057 2.002 1.964 1.075 1.014 0.968
$\textsc{REFRAG}_{8}$ 1.124 1.091 1.062 1.905 1.868 1.844 1.996 1.956 1.927 0.997 0.952 0.916
$\textsc{REFRAG}_{16}$ 1.157 1.114 1.076 1.925 1.882 1.853 2.016 1.971 1.938 1.034 0.976 0.931
$\textsc{REFRAG}_{32}$ 1.215 1.154 1.103 1.946 1.896 1.862 2.039 1.987 1.949 1.097 1.020 0.961
Table 2: Log-Perplexity on output tokens $x_{s+1:s+o}$ given different length of context. We use $s\in\{4096,8192,16384\}$ and $o=2048$ here. Bolding are based on comparing baselines excluding LLaMA-Full Context and LLaMA-32K since they are expected to be the best (ideally). The lower the better ($\mathbf{\downarrow}$).
Context Length =4096 Context Length=8192 Context Length=16384
Arxiv Book PG19 ProofPile Arxiv Book PG19 ProofPile Arxiv Book PG19 ProofPile $\mathbf{\downarrow}$
LLaMA-Full Context 6.751 6.956 6.829 6.701 9.675 9.069 8.963 9.401 9.043 8.913 8.848 8.989
LLaMA-32K 1.037 1.862 1.960 0.898 0.965 1.867 1.947 0.834 0.865 1.840 1.943 0.770
LLaMA-No Context 1.253 1.925 2.030 1.126 1.226 1.949 2.032 1.110 1.174 1.939 2.041 1.081
REPLUG 1.253 1.925 2.030 1.126 1.226 1.949 2.032 1.110 1.174 1.939 2.041 1.081
CEPE 1.085 1.856 1.959 0.945 1.032 1.878 1.958 0.904 0.960 1.864 1.966 0.863
$\textsc{REFRAG}_{8}$ 1.042 1.837 1.922 0.894 0.983 1.839 1.922 0.858 0.977 1.840 1.939 0.891
$\textsc{REFRAG}_{16}$ 1.058 1.847 1.934 0.910 0.994 1.845 1.932 0.871 0.942 1.840 1.945 0.850
$\textsc{REFRAG}_{32}$ 1.088 1.857 1.946 0.944 1.032 1.860 1.945 0.912 0.969 1.852 1.955 0.880
Figure 3: Log-Perplexity on $x_{s+1:s+o}$ under varying compression rates by selectively compressing different percentages of chunks. We compare three selection methods: RL (trained policy), Perplexity-desc (heuristic: lower perplexity), Perplexity-asc (heuristic: higher perplexity), and Random (random selection).
4.1 Ablation Study
Curriculum learning is essential for effective training in the reconstruction task. The reconstruction task, while intuitive, is particularly challenging when multiple chunks must be reconstructed. Table˜11 shows the performance of the reconstruction task with and without curriculum learning (i.e., reconstruction of $x_{1:s}$ from $s/k$ chunk embedding directly). The results indicate that curriculum learning is essential for the success of the reconstruction task.
Reconstruction task is essential for the model to learn the continual pre-training task. Table˜12 shows the performance of the continual pre-training task with and without initialization from the reconstruction task. The results indicate that pre-training on the reconstruction task is important for the success of continual pre-training.
Advantages of RL-based selective compression. Figure˜3 under various compression rates, achieved by varying the number of chunks to compress (i.e., adjusting $p$). Notably, a compression rate of $8$ can be obtained either by configuring $\textsc{REFRAG}_{16}$ to compress the appropriate number of chunks, or by employing $\textsc{REFRAG}_{8}$ with full compression, which is natively trained at a compression rate of $8$. This raises a natural question: does the former approach outperform the latter? Table˜13 demonstrates that $\textsc{REFRAG}_{16}$ with RL-based selective compression consistently outperforms $\textsc{REFRAG}_{8}$ across different datasets and context lengths. This finding is particularly surprising, as $\textsc{REFRAG}_{16}$ achieves a compression rate of $8$ without recomputing chunk embeddings, yet still surpasses the performance of $\textsc{REFRAG}_{8}$. These results further highlight the effectiveness of the RL-trained policy and underscore the practicality of dynamically adjusting the compression rate without compromising performance.
REFRAG trained under different compression rates. Figure˜10 illustrates the training trajectory of REFRAG under different compression rates in the continual pre-training task. We observe a performance regression as the compression rate increases; however, even at a compression rate of $32$, our model remains competitive (as shown in table˜1). In contrast, a compression rate of $64$ appears to be overly aggressive, resulting in diminished performance. These findings suggest a practical limit to the compression rate beyond which the model’s capability is significantly reduced.
Different combinations of encoder and decoder models for REFRAG. We employ LLaMA-2-7B and LLaMA-2-13B as decoders, and RoBERTa-Base and RoBERTa-Large as encoders, to investigate how model performance varies with different encoder and decoder sizes. Figure˜11 presents results for various encoder-decoder combinations. We observe that increasing the number of parameters in the decoder leads to a substantial reduction in loss, whereas enlarging the encoder yields only a modest improvement. This discrepancy may be attributed to the relatively minor increase in size from RoBERTa-Base to RoBERTa-Large compared to the substantial jump from 7B to 13B in the decoder. Additional results in figure˜12 indicate that a larger encoder may not always be advantageous when training with limited data in the continual pre-training setting. This observation aligns with previous findings by Li et al. (2024), which demonstrate that larger encoders in multi-modal models can negatively impact performance when data is scarce. To further validate our training approach on other decoder models, we conduct experiments with LLaMA-3.1-8B and LLaMA-3.2-3B. Table˜14 reports the performance of these models paired with RoBERTa-Base and RoBERTa-Large encoders on the Arxiv domain. Models trained with our recipe achieve performance comparable to the Full Context setting (i.e., without context compression). Moreover, increasing the context length continues to benefit our model, as evidenced by lower perplexity for a context length of $4096$ compared to $2048$.
5 Contextual Learning Applications
In this section, we investigate fine-tuning the model obtained from the pre-training stage to address various downstream tasks, including RAG, long document summarization, and multi-turn conversation with RAG. For each application, we curate an instruction-tuning dataset to facilitate model fine-tuning.
5.1 Retrieval Augmented Generation
Figure 4: RAG performance comparison under a strong retriever scenario (left) and a weak retriever scenario and a strong retriever scenario (right). REFRAG perform similarly to LLaMA model under the same retrieved passages (slightly better in a weaker retriever case) while outperform significantly under the same latency.
Training dataset. We follow the work of Lin et al. (2024) and use a combination of question answering datasets from 5 domains to fine-tune our model, which contains 1.1 million data points. Dialogue: OpenAssistant Conversations Dataset. Open-Domain QA: CommonsenseQA, MathQA, Web Questions, Wiki Question Answering, Yahoo! Answers QA, FreebaseQA, MS MARCO. Reading Comprehension: Discrete Reasoning Over Paragraphs, PubMedQA, QuaRel, SQuADv2. Chain-of-thought Reasoning: Algebra QA with Rationales, Explanations for CommonsenseQ, Grade School Math 8K, MathQA, StrategyQA.
Evaluation dataset. We hold out 5% of the data for each dataset in the training dataset for evaluation. Additionally, we use the datasets that are commonly used in RAG literature (Izacard et al., 2023b; Lin et al., 2024), including MMLU (Hendrycks et al., 2021), BoolQ (Clark et al., 2019), SIQA (Sap et al., 2019), PIQA (Bisk et al., 2020), and Knowledge Intensive Language Tasks (KILT) (Petroni et al., 2020) (including HellaSwag, Winogrande, TQA, FEVER, NQ). We evaluate our performance on 2 settings: 1) Strong Retriever: In this setting we use a strong retriever and retrieve the K-nearest neighbors to answer the question; 2) Weak Retriever: In this setting we retrieve 200 passages and pick random K passages to answer the question. The weak retriever setting closely resembles real-world systems, as RAG retrieval systems often suffer from error accumulation across subsystems. A table summarizing the evaluation metrics for each dataset is included in table˜7.
Retriever and retrieval corpus. We follow the work of Lin et al. (2024) to use Wikipedia dumps and CommonCrawl dumps to create a retrieval corpus with 400 million passages. Each passage contains less than 200 words. We use the DRAGON+ model Lin et al. (2023) as our retriever and use the implementation of Izacard et al. (2023a) to retrieve the K-nearest neighbors as the retrieved passages for each question.
Result analysis. Table˜3 shows the performance of different baselines under short and long contexts (i.e., varying number of retrieved passages) [^3] Note that the implementation of our exact match is stricter than other works. We follow the work of Lin et al. (2024) to use the stricter version and hence the reported numbers are lower in general.. (1/# tokens) is inverse for the number of tokens in the decoder model. This is used as a metric to gauge the latency of the model (the higher, the lower latency). $\textsc{LLaMA}_{\text{FT}}$ is the original LLaMA-2-7B model that is fine-tuned on the same RAG dataset used to train our model. We compare the performance under both the short context and the long context scenarios. For the short context, we use 1 passage for $\textsc{LLaMA}_{\text{FT}}$ and use 8 passages for all our models. The baseline of $\textsc{REFRAG}_{8}$ will have the same latency as the $\textsc{LLaMA}_{\text{FT}}$ model. However, due to the compression, we are able to have more context information and hence achieve better performance. Surprisingly, $\textsc{REFRAG}_{16}$ and $\textsc{REFRAG}_{32}$ both outperform the $\textsc{LLaMA}_{\text{FT}}$ model despite having $2\times$ and $4\times$ fewer tokens in the decoder (i.e., lower latency). The same result occurs in long context scenarios. Our model has even higher performance gains in multi-choice tasks. Table˜15 shows the performance of our model under different numbers of passages. The result suggests that most tasks still benefit from more passages in our model. Figure˜4 shows the performance averaged over all 16 tasks in table˜3 for both strong retriever and weak retriever setting. The result demonstrates that under the same number of retrieved passages, we are able to match the performance of LLaMA in the strong retriever setting and even outperform LLaMA under the weak retriever setting. This is because our model enables larger context and hence enables extract more useful information when the retrieved passages are less relevant. Under equivalent latency constraints, REFRAG consistently outperform LLaMA on both settings as the saved context can be reinvested to include additional information within the same latency budget.
Figure˜4 compares the performance of REFRAG and the LLaMA model under two conditions: 1) an equal number of retrieved passages, and 2) equal latency, for both strong and weak retriever settings. With a strong retriever and a maximum of 10 passages, REFRAG matches LLaMA’s performance while achieving a $5.26\times$ speedup in TTFT. At equal latency (8 passages for REFRAG vs. 1 for LLaMA), REFRAG attains a $1.22\%$ average improvement across 16 RAG tasks. With a weak retriever setting, at 10 passages, REFRAG improves performance by $0.71\%$ and accelerates TTFT by $5.26\times$ compared to LLaMA. At equal latency (8 passages for REFRAG vs. 1 for LLaMA), REFRAG achieves a $1.93\%$ average gain over 16 RAG tasks.
Table 3: Comparison of model performance of different models with different number of retrieved passages for RAG under the strong retriever scenario.
Generation NQ FEVER TQA WebQA FreebaseQA GSM8K StrategyQA BoolQ $\mathbf{\uparrow}$ (1/ # tokens)
Short context with the same latency
$\textsc{LLaMA}_{\text{FT}}$ + 1 passage 23.96 62.04 9.64 37.33 75.18 7.38 64.44 29.24 $1\times$
$\textsc{REFRAG}_{8}$+ 8 passages 22.96 66.59 13.05 38.67 73.46 7.38 75.56 3.30 $1\times$
$\textsc{REFRAG}_{16}$+ 8 passages 22.94 62.88 12.97 42.67 71.50 9.40 71.11 5.87 $2\times$
$\textsc{REFRAG}_{32}$+ 8 passages 22.11 64.24 12.57 41.33 71.74 12.75 73.33 1.99 $4\times$
Long context
$\textsc{LLaMA}_{\text{FT}}$ + 10 passages 26.08 65.44 9.68 40.00 76.17 6.71 68.89 30.00 $1\times$
CEPED +80 passages 0.03 65.68 0.01 0.00 0.00 0.00 0.00 57.80
REPLUG +80 passages - - - - - - 64.44 -
LLaMA-32K +80 passages 1.24 0.14 0.52 10.67 9.83 0.00 0.00 0.03
$\textsc{REFRAG}_{8}$ +80 passages 24.15 68.83 13.06 37.33 74.20 7.38 71.11 7.03 $1\times$
$\textsc{REFRAG}_{16}$ +80 passages 23.30 66.01 12.65 38.67 75.43 12.08 73.33 12.23 $2\times$
$\textsc{REFRAG}_{32}$ +80 passages 23.02 68.48 12.14 38.67 71.74 9.40 68.89 6.42 $4\times$
Multi-Choice MMLU CommonsenseQA MathQA ECQA HellaSwag SIQA PIQA Winogrande $\mathbf{\uparrow}$
Short context with the same latency
$\textsc{LLaMA}_{\text{FT}}$ + 1 context 50.23 85.05 99.50 84.77 41.80 68.12 67.36 55.64 $1\times$
$\textsc{REFRAG}_{8}$ + 8 passages 50.29 92.27 99.66 94.70 45.23 68.94 71.38 57.70 $1\times$
$\textsc{REFRAG}_{16}$ + 8 passages 49.84 89.18 99.66 98.01 39.33 68.42 70.29 56.67 $2\times$
$\textsc{REFRAG}_{32}$ + 8 passages 49.51 91.75 99.50 97.35 42.86 68.17 68.34 56.75 $4\times$
Long context
$\textsc{LLaMA}_{\text{FT}}$ + 10 passages 48.66 82.99 68.46 84.11 41.77 67.45 68.01 53.91 $1\times$
CEPED +80 passages 26.26 26.29 23.66 24.50 24.95 32.86 48.53 44.51
REPLUG +80 passages - 78.35 - 76.16 - 65.51 - -
LLaMA-32K +80 passages 22.21 16.49 19.80 16.56 23.76 24.16 34.17 48.86
$\textsc{REFRAG}_{8}$ +80 passages 50.42 92.27 99.66 97.35 44.61 68.22 69.37 57.54 $1\times$
$\textsc{REFRAG}_{16}$ +80 passages 50.88 89.69 99.66 96.69 38.50 68.47 70.89 56.99 $2\times$
$\textsc{REFRAG}_{32}$ +80 passages 49.77 90.72 99.50 98.01 43.37 68.47 69.04 56.99 $4\times$
- means the corresponding model has out-of-memory error.
5.2 Multi-Turn Conversation
We use three different knowledge-intensive multi-turn conversation datasets: TopiOCQA (Adlakha et al., 2022), ORConvQA (Qu et al., 2020), and QReCC (Anantha et al., 2021). For each conversation turn, we retrieve $K$ passages using the same retriever and retrieval corpus as described in section˜5.1.
Result analysis. Table˜4 presents results across varying numbers of conversational turns and retrieved passages. Our model outperforms $\textsc{LLaMA}_{\text{FT}}$ on two out of three datasets in the 5-passage setting, and on all three datasets in the 10-passage setting. This improvement is attributable to the limited 4k-token context window of $\textsc{LLaMA}_{\text{FT}}$, which necessitates truncating portions of the conversational history in longer contexts, resulting in the loss of crucial information required to answer subsequent questions. In contrast, our model, trained on the same LLaMA model without extending its effective positional encoding, maintains robust performance even with a large number of passages, owing to the benefits of our compression approach. Table˜5 further reports the performance of different models under varying numbers of passages, with our model consistently achieving superior results on two out of three datasets for the reasons outlined above.
Table 4: Performance on multi-turn RAG tasks for # Passages = 5 and # Passages = 10.
# Turns ($\geq$) ORConvQA QReCC TopiOCQA $\mathbf{\uparrow}$
# Passages = 5
$\textsc{LLaMA}_{\text{FT}}$ 2 20.73 18.72 26.98
$\textsc{REFRAG}_{8}$ 2 21.17 17.73 28.04
$\textsc{REFRAG}_{16}$ 2 20.19 17.30 27.89
$\textsc{REFRAG}_{32}$ 2 19.70 17.35 28.67
$\textsc{LLaMA}_{\text{FT}}$ 4 20.33 16.42 23.50
$\textsc{REFRAG}_{8}$ 4 22.78 15.61 26.93
$\textsc{REFRAG}_{16}$ 4 21.94 15.27 27.03
$\textsc{REFRAG}_{32}$ 4 21.68 15.45 26.45
$\textsc{LLaMA}_{\text{FT}}$ 6 20.76 11.92 23.10
$\textsc{REFRAG}_{8}$ 6 23.11 10.88 25.37
$\textsc{REFRAG}_{16}$ 6 21.69 10.75 26.17
$\textsc{REFRAG}_{32}$ 6 21.19 10.69 25.51
# Turns ($\geq$) ORConvQA QReCC TopiOCQA $\mathbf{\uparrow}$
# Passages = 10
$\textsc{LLaMA}_{\text{FT}}$ 2 16.52 17.31 23.02
$\textsc{REFRAG}_{8}$ 2 21.15 17.92 27.97
$\textsc{REFRAG}_{16}$ 2 20.79 17.37 28.45
$\textsc{REFRAG}_{32}$ 2 19.67 17.16 28.31
$\textsc{LLaMA}_{\text{FT}}$ 4 16.90 14.69 20.23
$\textsc{REFRAG}_{8}$ 4 22.63 15.68 25.95
$\textsc{REFRAG}_{16}$ 4 21.84 15.21 26.12
$\textsc{REFRAG}_{32}$ 4 21.75 15.33 25.77
$\textsc{LLaMA}_{\text{FT}}$ 6 14.44 10.72 19.52
$\textsc{REFRAG}_{8}$ 6 20.59 11.00 25.16
$\textsc{REFRAG}_{16}$ 6 21.05 10.50 24.96
$\textsc{REFRAG}_{32}$ 6 21.67 10.79 25.23
Table 5: Performance on multi-turn RAG tasks with different number of passages.
REFRAG $\textsc{LLaMA}_{\text{FT}}$
# Passages ORConvQA QReCC TopiOCQA $\mathbf{\uparrow}$ ORConvQA QReCC TopiOCQA $\mathbf{\uparrow}$
0 19.27 15.32 28.19 19.16 15.49 28.22
5 20.18 17.37 28.24 19.65 18.71 27.08
8 20.52 17.60 28.17 16.87 18.05 25.36
10 19.67 17.41 27.62 15.72 17.42 23.60
6 Related Works
Retrieval-Augmented Language Modeling.
Recent research has extensively investigated novel model architectures to improve retrieval-augmented generation. Guu et al. (2020) introduced pre-training for retrieval-augmented masked language models. Building on this, Borgeaud et al. (2022) proposed a new architecture and pre-training paradigm for generative LLMs, leveraging cross-attention and end-to-end pre-training with retrieval from a trillion-token data store, achieving strong performance. Subsequent work by Shi et al. (2024) and Lin et al. (2024) focused on fine-tuning existing LLMs by prepending retrieved passages to prompts and employing ensemble methods for response generation. Additionally, Izacard and Grave (2021) introduced fusion-in-decoder, which uses an encoder to process each passage in parallel and concatenates the hidden states for generation via a decoder. This approach accelerates attention computation by removing cross-document attention, but does not apply compression in the decoder, which could further reduce latency.
Efficient Long-Context LLMs.
Recent research has investigated various strategies to reduce memory usage and accelerate latency in long-context generation for LLMs. Choromanski et al. (2021) introduced compressed attention, reducing attention complexity from quadratic to linear; however, this method does not address memory requirements. It is complementary to our approach and can be integrated to further improve latency. StreamingLLM(Xiao et al., 2024) proposed attention sinks to decrease KV cache memory for long-context generation, though this does not reduce latency during the pre-filling stage. CEPE (Yen et al., 2024) employs cross-attention to token embeddings from context tokens, reducing both KV cache memory and attention computations. However, CEPE is limited to prefix context applications, as it disrupts the causal structure of the context, making it unsuitable for tasks such as multi-turn RAG or summarization. Additionally, CEPE does not utilize token compression, resulting in similar or even increased decoding latency. Concurrently with our work, Dai et al. (2025) proposed PCC, an embedding-based memory mechanism that summarizes past context into compact vectors, enabling retrieval of salient information during subsequent processing. Like CEPE, PCC is limited to prefix context applications and does not support arbitrary folding or expansion of contexts at any position. Interestingly, Kuratov et al. (2025) investigated the capacity of LLMs to encode long contexts into a single embedding, demonstrating minimal information loss for sequences up to 1500 tokens. Their work examines the extent to which information can be compressed into a single embedding, offering a complementary perspective to REFRAG, which is designed for decoding from multiple compact embeddings within the standard decoder architecture.
Compressive transformer.
Rae et al. (2020) first introduced the compressive transformer, which compresses the KV cache to reduce memory requirements for long-context applications. However, this approach only decreases KV cache memory usage, does not improve time-to-first-token latency, and requires training the model from scratch. Yoshida et al. (2021) extended this idea by employing recursive context compression, generating a summary hidden state for each chunk to inform the next chunk’s computation. The recursive nature, however, prevents pre-computation and reuse of chunk embeddings, and does not reduce decoding latency. Chevalier et al. (2023) proposed recursive compression for documents, using compressed embeddings for prediction, similar to our method. However, their sequential compression process results in high latency when the summary vector is not cached, and their approach only supports applications where the summary token is restricted to the prefix of the language model (e.g., RAG), limiting applicability. In contrast, our work is the first to enable pre-computation of chunk embeddings and their use at arbitrary positions within the prompt, supporting diverse applications such as RAG and multi-turn conversation. Furthermore, our method learns where to apply compression, allowing for adaptive compression rates at inference time without recomputing chunk embeddings.
Prompt compression.
Prompt compression seeks to reduce input token length to lower latency and cost while maintaining task performance. A prominent approach is LLMLingua(Jiang et al., 2023),which employs coarse-to-fine, budget-controlled compression with token-level iterative refinement, achieving high compression ratios with minimal performance loss. LongLLMLingua (Jiang et al., 2024) extends this method to long-context scenarios, demonstrating significant cost and end-to-end speed improvements. Complementary approaches rank or prune context by estimated informativeness, e.g., Selective Context uses self-information to drop low-value tokens, and sentence-level methods learn context-aware encoders for question-specific compression and faster inference Li et al. (2023); Liskavets et al. (2024). These approaches are complementary to our work and can be integrated to further reduce the latency of REFRAG.
7 Conclusion
In this work, we introduced REFRAG, a novel and efficient decoding framework tailored for RAG applications. By leveraging the inherent sparsity and block-diagonal attention patterns present in RAG contexts, REFRAG compresses, senses, and expands context representations to significantly reduce both memory usage and inference latency, particularly the TTFT. Extensive experiments across a range of long-context applications, including RAG, multi-turn conversations, and long document summarization, demonstrate that REFRAG achieves up to $30.85\times$ TTFT acceleration ($3.75\times$ over previous state-of-the-art methods) without any loss in perplexity or downstream accuracy. Our results highlight the importance of specialized treatment for RAG-based systems and open new directions for efficient large-context LLM inference. We believe that REFRAG provides a practical and scalable solution for deploying LLMs in latency-sensitive, knowledge-intensive applications.
8 Acknowledgements
We thank for Jason Chen, Yao Liu, Norman Huang, Xueyuan Su, Pranesh Srinivasan, Avinash Atreya, Riham Mansour, Jeremy Teboul for insightful discussions and support.